rellie invests $12,000.a part of it earns 10% annual interest and the rest earns 9% annual interest.after a year, he earned $1,130 from his investments.how much is invested at each rate?

Answer by mananth(16949)

let him invest $ x at 10%
so 12000-x will be invested at 9%
interest earned on both =$1130
..
0.1x+0.09(12000-x)=1130
0.1x+1080-0.09x=1130
0.01x=1130-1080
0.01x=50
x=50/0.01
x= 5000
$5000 invested @10%
$7000 invested @9%